1. Find the value of R, ignoring early effect for DC bias calculation

2. Find the small signal parameters of the amplifier

3. Draw the small signal equivalent circuit

4. Calculate Avo, Ri, and Ro

15 V R + 0.1 uF V;GH VCE Vo R 3 ΚΩ

VCE = 8.2 V

Rsig = 0

RL = infinity

Va = 320 V

Beta = 100
Answer #1

us DC. ANALYSIS 15 v sk Let Assume Given, Vee = 8-2 V fc R + BE100 Ic , Ip and IB all are in mA. VCE VA = 320V If .SE R ww 32) * А Apply Kun through B Pay : Э «, — TuR — З Те – о - ) 15 2 — 2Ге с - Тg x R 2. ә) 7. — 8 = 16x R ; *2, к. R= 1.4 x(3 +na a 1 - 16.kn Ang 13) Collector current , lc = B 1+ß TE - 1oo 101 X 6.8 mĄ 3 Gas Ic= 2.244 mA A small signal parameter(s) :(4) circuits 3) small signal signal Equivalent B Ib С Vj re ro R3 § R •B16 Ę L Rj Vo 3 Kg Ro where, Va Doa = 320 Ka 2.244 1cGain ( Anolio • Apply Kch at Emitter Node: Vogue Vo You Vj ß Ib = t + Do ra and Vg - Vo I b (2) ra » 3 ra using eqn . in egr